A flag pole 24 feet high is 10 feet away from the base of its anchor cable. How long is the anchor cable?
Answer by unlockmath(1688)

Hello,
Since this forms a right triangle we can use this formula:
A^2 + B^2 = C^2 C is going to be the hypotenuse (diagonal line) or in this case the anchor cable. Now we can plug in the numbers
24^2 + 10^2 = c^2 Do the calculations will give us:
576 + 100 = c^2 rewritten will be
676 = c^2 Now square root both sides will be:
26 = C There we go the length is 26 feet.
RJ Toftness
Check out my book at www.math-unlock.com